OpenStudy (anonymous):

The two rectangles below are similar first one (bottom is 12 cm, and right side is 5 cm) Second one(Bottom (2x + 8)cm) right side x cm) Its not C because thats the answer i used and i got it wrong

OpenStudy (campbell_st):

you need to use the ratio or corresponding sides in similar rectangles. Which will be in the same ratio \[\frac{2x + 8}{12} = \frac{x}{5}\] cross multiply and then solve for x